The 𝕓𝕓𝕠𝕝𝕕 symbol font

Alan Jeffrey

April 1994

Introduction

This document describes the 𝕓𝕓𝕠𝕝𝕕 math alphabet. This is an open (or ‘blackboard bold’) geometric sans serif, intended mainly for use in mathematics, but which may prove useful for decorative work.

To use the fonts in L^{A}TEX 2ε, you select the bbold package, and then use the \mathbb command to get blackboard bold mathematics, \textbb for text, or \bbfamily for longer text. The full font is:

´ 0  ´ 1  ´ 2  ´ 3  ´ 4  ´ 5  ´ 6  ´ 7
´ 00x  ℾ  Δ  Θ  Λ  Ξ  ℿ  ⅀  Υ ˝ 0x ´ 01x  Φ  Ψ  Ω  α  β  ℽ  δ  ε
´ 02x  ζ  η  θ  ι  κ  λ  μ  ν ˝ 1x ´ 03x  ξ  ℼ  ρ  σ  τ  υ  ϕ  χ
´ 04x  ψ  !  ”  #  $  %  &  ’ ˝ 2x ´ 05x  ⦅  ⦆  *  +  ,  -  .  /
´ 06x  𝟘  𝟙  𝟚  𝟛  𝟜  𝟝  𝟞  𝟟 ˝ 3x ´ 07x  𝟠  𝟡  :  ;  <  ·  >  ?
´ 10x  @  𝔸  𝔹  ℂ  𝔻  𝔼  𝔽  𝔾 ˝ 4x ´ 11x  ℍ  𝕀  𝕁  𝕂  𝕃  𝕄  ℕ  𝕆
´ 12x  ℙ  ℚ  ℝ  𝕊  𝕋  𝕌  𝕍  𝕎 ˝ 5x ´ 13x  𝕏  𝕐  ℤ  ⟦  \  ⟧  ⟪  ⟫
´ 14x  ‘  𝕒  𝕓  𝕔  𝕕  𝕖  𝕗  𝕘 ˝ 6x ´ 15x  𝕙  𝕚  𝕛  𝕜  𝕝  𝕞  𝕟  𝕠
´ 16x  𝕡  𝕢  𝕣  𝕤  𝕥  𝕦  𝕧  𝕨 ˝ 7x ´ 17x  𝕩  𝕪  𝕫  –  ⫾  —  “  ω

           ˝ 8    ˝ 9    ˝ A    ˝ B    ˝ C    ˝ D    ˝ E    ˝ F

𝕋𝕙𝕖 𝕓𝕓𝕠𝕝𝕕 𝕗𝕠𝕟𝕥 𝕗𝕒𝕞𝕚𝕝𝕪.

This was the first full font I implemented, and has a number of dubious design features. It is very geometric, and the stress does not conform to that of most Latin fonts—this was an attempt to imitate real blackboard handwriting. The difference in widths between ‘𝕒’ and ‘𝕓’ is very noticable. The lip of the ‘𝕣’ is horrible. However, the uppercase latin letters look passable, and they’re almost the only ones anyone ever uses.

Legal rubbish

This package is copyright \begin{aligned}& c \\ ◯ \end{aligned} 1989–1994 Alan Jeffrey. All rights are reserved. The moral right of the author has been asserted.

You are not allowed to take money for the distribution or use of this file except for a nominal charge for copying, etc.

Redistribution of unchanged files is allowed provided that the whole package is distributed.